Problem 5-3 Calculating Present Values [LO2]

For each of the following, compute the present value (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.):

  

Present Value Years Interest Rate Future Value
14 8 % $15,551
5 14 52,557
30 15 887,073
35 8 551,164
Answer #1

1.Information provided:

Future value= $15,551

Time= 14 years

Interest rate= 8%

The present value is calculated by entering the below in a financial calculator:

FV= 15,551

N= 14

I/Y= 8

Press the CPT key and PV to compute the present value.

The value obtained is 5,294.51.

Therefore, the present value is $5,294.51.

2. Information provided:

Future value= $52,557

Time= 5 years

Interest rate= 14%

The present value is calculated by entering the below in a financial calculator:

FV= 15,551

N= 14

I/Y= 8

Press the CPT key and PV to compute the present value.

The value obtained is 5,294.51.

Therefore, the present value is $5,294.51.

3. Information provided:

Future value= $887,073

Time= 30 years

Interest rate= 15%

The present value is calculated by entering the below in a financial calculator:

FV= 887,073

N= 30

I/Y= 15

Press the CPT key and PV to compute the present value.

The value obtained is 13,397.51.

Therefore, the present value is $13,397.51.

4. Information provided:

Future value= $551,164

Time= 30 years

Interest rate= 15%

The present value is calculated by entering the below in a financial calculator:

FV= 887,073

N= 35

I/Y= 8

Press the CPT key and PV to compute the present value.

The value obtained is 8,324.26.

Therefore, the present value is $8,324.26.

In case of any query, kindly comment on the solution.